OpenStudy (anonymous):

cot(x)sec^4(x)=cot(x)+2tan(x)+tan^x please do this step by step so i can get a good grasp on what is happening

OpenStudy (anonymous):

USE: sec^2(x) = 1+tan^2(x)

OpenStudy (anonymous):

how did you get that.

OpenStudy (anonymous):

Well, sec^2(x) = 1+ tan^2(x) is an identity

OpenStudy (anonymous):

like sin^2(x) + cos^2(x)=1

OpenStudy (anonymous):

from the defination of trigonometric ratios, we have 

OpenStudy (anonymous):

now, sec^2(x) = h^2/b^2 = (p^2 + b^2)/b^2 = p^2/b^2 +1 = tan^2(x) + 1

OpenStudy (anonymous):

i just need to prove the identity

OpenStudy (anonymous):

Now, you need to use sec^2(x) = 1 + tan^2(x) to Prove cot(x)sec^4(x)=cot(x)+2tan(x)+tan^3(x)